Fighter Jet pilots can be subjected to quite large forces during flight maneuvers. A 80 kg pilot is flying in a swooping arc, down and then up, with a radius of 500m and a speed of 140 m/s. What is the normal force of the seat on the pilot when she is at the lowest point of the arc?
1
Expert's answer
2020-10-30T13:17:10-0400

"N=mg+m\frac{v^2}{r}\\N=(80)(9.8)+80\frac{140^2}{500}=3920\ N"